# Simulating stochastic population dynamics: The Linear Noise Approximation can capture non-linear phenomena

### v3

@@ 3.1 The pcLNA algorithm for Hopf bifurcation systems
We now wish to define the phase correction map $G$ for a reaction network with macroscopic RRE as in (6) satisfying conditions (H.I), (H.II). Consider some value $\alpha$ of the bifurcation parameter that is near the critical $\alpha_{p}$ and suppose that ${\bm{x}}_{\alpha}$ is an equilibrium for this value of the bifurcation parameter, i.e. ${\bm{F}}({\bm{x}}_{\alpha},\alpha)={\bm{0}}$. Consider also a deterministic trajectory $\gamma=\{{\bm{x}}(t):t\geq 0\}$ near ${\bm{x}}_{\alpha}$ We wish to define the phase map ${G}:\mathbb{R}^{n}\to\gamma$ for any stochastic state ${\bm{X}}\in\mathbb{R}^{n}$.
First, we find the $\bm{u}$} coordinates in (28). We identify the conjugate pair of eigenvectors, ${\bm{u}}_{1}\pm i{\bm{u}}_{2}$, corresponding to the complex eigenvalues of the Jacobian $\bm{J}({\bm{x}}_{\alpha},\alpha)$ The eigenspace corresponding to these complex eigenvalues is spanned by $\{{\bm{u}}_{1},{\bm{u}}_{2}\}$, and while the vectors ${\bm{u}}_{1},{\bm{u}}_{2}$ are not necessarily orthonormal, we can easily find an orthonormal basis $\{\hat{{\bm{u}}}_{1},\hat{{\bm{u}}}_{2}\}$ of this eigenspace using Gram-Schmidt orthonormalisation or a similar method. Letting ${\bm{R}}_{1}=[\hat{{\bm{u}}}_{1}\;\hat{{\bm{u}}}_{2}]$ we define the map $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}^{2}$, ${\bm{u}}=\text{proj}({\bm{x}})={\bm{R}}_{1}^{T}({\bm{x}}-{\bm{x}}_{\alpha})$. We use the projection map to derive the $\bm{u}$}-coordinates for the deterministic trajectory $\gamma$, $\{{\bm{u}}(t):t\geq 0\}$ with ${\bm{u}}(t)=\text{proj}({\bm{x}}(t))$, and for any stochastic state $\bm{X}$}, ${\bm{U}}=\text{proj}({\bm{X}})$.
As discussed in section 2.5.6, to control phase drifts we need to control the deviations of the stochastic trajectory in the $\bm{u}$} coordinates. We therefore wish to minimise the distance of any stochastic states $\bm{U}$} to $\{{\bm{u}}(t):t\geq 0\}$. We define the map $\mathcal{M}\to\mathbb{R}^{2}\in(0,\infty)$, where $s=\mathcal{M}({\bm{U}})=\operatorname*{argmin}_{t\geq 0}\lVert{\bm{U}}-{\bm{u}}(t)\rVert$, that gives the value of time $t$ minimising the distance between $\bm{U}$} and $\gamma$. Finally, the map ${\bm{x}}:(0,\infty)\to\gamma$, ${\bm{x}}={\bm{x}}(s)\in\gamma$ simply lifts back the state ${\bm{u}}(s)$ to its original coordinates. The phase correction map $G$ is the convolution $G={\bm{x}}\circ\mathcal{M}\circ\text{proj}$, ${\bm{x}}(s)=G({\bm{X}})={\bm{x}}(\mathcal{M}(\text{proj}({\bm{X}}))$. The perturbation $\bm{\kappa}(t)=\sqrt{\Omega}({\bm{X}}(t)-G({\bm{X}}(t)))$ has minimised $\bm{u}$} coordinate (see Figure 6).
Having defined the map G{G}, we provide the pcLNA algorithm for Hopf bifurcation systems in Algorithm 3.
@@
a.. Initial conditions: $s_{0}=0=t_{0}$, ${\bm{x}}(0)$, ${\bm{X}}(0)$, Parameters: $\Delta t$, $\Omega$.
b. System satisfying (H.I), (H.II), with solution $\gamma=\{{\bm{x}}(t):t\geq 0\}$ and the corresponding solutions $\bm{C}$} and $\bm{D}$} of (12), (13), respectively.
c. The projection map $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}^{2}$, ${\bm{u}}=\text{proj}({\bm{x}})=\mathbb{R}_{1}^{T}({\bm{x}}-{\bm{x}}_{\alpha})$ where ${\bm{R}}_{1}=[\hat{{\bm{u}}}_{1}\;\hat{{\bm{u}}}_{2}]$ with $\hat{{\bm{u}}}_{1}$, $\hat{{\bm{u}}}_{2}$ an eigenbasis corresponding to the conjugate pair of complex eigenvalues of $\bm{J}({\bm{x}}_{\alpha},\alpha)$, and ${\bm{x}}_{\alpha}$ such that ${\bm{F}}({\bm{x}}_{\alpha},\alpha)={\bm{0}}$.
Steps: Compute $\{{\bm{u}}(t):t\geq 0\}$, where ${\bm{u}}(t)=\text{proj}({\bm{x}}(t))$ and $\bm{\kappa}(s_{0})=\Omega^{1/2}({\bm{X}}(0)-{\bm{x}}(0))$.
@@ 4 Bi-stable systems
We now consider systems that present bi-stability. We focus on the type of bi-stability arising in the systems presented in [2], [19] and [12]. We wish to derive stochastic trajectories $\{{\bm{X}}(t_{i}):i=0,1,2,\dots\}$, with ${\bm{X}}(t_{i})\in\mathbb{R}^{n}$, for a reaction network with RRE limit (as $\Omega\to\infty$) satisfying (6) and Jacobian, $\bm{J}({\bm{x}},\alpha)$, as in (7) that satisfy the conditions:
@@
    for $\alpha$ near $\alpha_{p}\in\mathbb{R}$ there exist ${\bm{x}}_{1}^{*}\in\mathbb{R}^{n}$ such that ${\bm{F}}({\bm{x}}_{1}^{*},\alpha)={\bm{0}}$ and all the eigenvalues of the Jacobian matrix $\bm{J}({\bm{x}}_{1}^{*},\alpha)$ are negative.
The last condition (B.III) ensures the presence of a stable equilibrium point for all values of $\alpha$ near $\alpha_{p}\in\mathbb{R}$. The conditions (B.I-II) prescribe the presence of a non-hyperbolic equilibrium, ${\bm{x}}_{\alpha_{p}}$, and the existence of a parameter dependent, one-dimensional, attracting manifold near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$ that is l.t.e. to a system of the form as in (15). Furthermore, the same conditions imply that, near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$, the system is l.t.e. to the Generic Fold bifurcation system. For the Generic Fold bifurcation system, the non-hyperbolic equilibrium present at $\alpha_{p}$ disappears for $\alpha<\alpha_{p}$ while two equilibrium points, one stable and one unstable, emerge for $\alpha>\alpha_{p}$.
Therefore, a system satisfying the conditions (B.I-III) presents two stable and one unstable equilibrium points for $\alpha>\alpha_{p}$. That is, we arrive to the following result.
@@
For a system as in (6) satisfying the conditions (B.I)-(B.III), with value of $\alpha$ such that $\alpha-\alpha_{p}$ is positive and and sufficiently small, there exist ${\bm{x}}_{1}^{*},{\bm{x}}_{\alpha},{\bm{x}}_{2}^{*}\in\mathbb{R}^{n}$ with ${\bm{F}}({\bm{x}}_{1}^{*},\alpha)={\bm{F}}({\bm{x}}_{\alpha},\alpha)={\bm{F}}({\bm{x}}_{2}^{*},\alpha)={\bm{0}}$ and all the eigenvalues of the Jacobian matrices $\bm{J}({\bm{x}}_{1}^{*},\alpha)$ and $\bm{J}({\bm{x}}_{2}^{*},\alpha)$ are negative (linear stability), while $\bm{J}({\bm{x}}_{\alpha},\alpha)$ has one positive eigenvalue (instability) and the rest are negative.
As mentioned above, we observe this type of bi-stability for the systems in [2], [19] and [12].
@@ 4.1 The pcLNA algorithm for bi-stable systems
Following the pcLNA principle, we construct a phase correction map $G$ to control phase drifts for systems satisfying these conditions. As discussed in section 2.3.1, the LNA is long-time accurate when the system is in the vicinity of a stable equilibrium. This is the case for $\alpha\leq\alpha_{p}$. For the case where there is bi-stability, that is when $\alpha>\alpha_{p}$, the standard LNA that uses a single deterministic solution to simulate a trajectory will produce trajectories centered around the “nearest” equilibrium and therefore it will fail when there is a significant likelihood that a stochastic trajectory produced by SSA escapes from the nearest equilibrium (see Figure 2(D) and SI Figure S19). The main focus of phase correction is therefore to allow for this event. At the very least, more than one deterministic trajectories will need to be used.
There is a simple way to predict the equilibrium that the current state is expected to converge. If ${\bm{R}}_{1}={\bm{u}}$ is the eigenvector corresponding to the eigenvalue of $\bm{J}({\bm{x}}_{\alpha},\alpha)$ which is zero when $\alpha=\alpha_{p}$, then we define the projection map $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}$, giving the $u$-coordinate $u=\text{proj}({\bm{x}})={\bm{R}}_{1}({\bm{x}}-{\bm{x}}_{\alpha})$. The $u$-coordinates of the two stable equilibria, $u^{*}_{1}=\text{proj}({\bm{x}}^{*}_{1})$ and $u^{*}_{2}=\text{proj}({\bm{x}}^{*}_{2})$, have different sign, say $u^{*}_{1}<0<u^{*}_{2}$. Furthermore, near ${\bm{x}}_{\alpha}$, the eigenvector $\bm{u}$} defines a separatrix, that is the trajectories with initial condition that has negative (positive) $u$-coordinate converge to ${\bm{x}}^{*}_{1}$ (${\bm{x}}^{*}_{2}$).
The phase map $G({\bm{X}})$ for a state $\bm{X}$} of the stochastic trajectory is chosen to control perturbations in the $u$-coordinate. For this, it is essential that $G({\bm{X}})$ and $\bm{X}$} have $u$-coordinate of the same sign. We consider a phase map that uses two trajectories $\gamma_{1}=\{{\bm{x}}^{(1)}(t):t\geq 0\}$ and $\gamma_{2}=\{{\bm{x}}^{(2)}(t):t\geq 0\}$, the first converging to ${\bm{x}}^{*}_{1}$ and the second to ${\bm{x}}^{*}_{2}$ and their corresponding $u$-coordinates are $\{u^{(j)}(t):t\geq 0\}$ with $u^{(j)}(t)=\text{proj}({\bm{x}}^{(j)}(t))$, $j=1,2$. Let $U=\text{proj}({\bm{X}})$. The phase map is then $G({\bm{X}})={\bm{x}}_{j}(s)$, where $j\in\{1,2\}$ satisfies $u_{j}^{*}\,U>0$ (i.e. $\bm{U}$} and $u_{j}^{*}$ have the same sign) and $s=\operatorname*{argmin}_{t\geq 0}\|U-u_{j}(t)\|$. This phase map ensures that when the trajectory jumps from one side of the separatrix to the other, then the map also switches side. Using only two trajectories is the most minimal option; one may consider using more trajectories or using a new trajectory every time the state jumps to a different side of the separatrix, but this will be computationally more costly. In Algorithm 4, we describe the associated pcLNA simulation algorithm. In section 5, we show that this simple phase map is sufficient to achieve high levels of accuracy.
@@
c. The projection map $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}$, $u=\text{proj}({\bm{x}})=\mathbb{R}_{1}^{T}({\bm{x}}-{\bm{x}}_{\alpha})$ where ${\bm{R}}_{1}={\bm{u}}$ with $\bm{u}$} a normalised eigenvector corresponding to the eigenvalue of $\bm{J}({\bm{x}}_{\alpha},\alpha)$ that is equal to zero for $\alpha=\alpha_{p}$.
Steps: Compute $\{u^{(j)}(t):t\geq 0\}$ where $u^{(j)}(t)=\text{proj}({\bm{x}}^{(j)}(t))$ and $u_{j}^{*}=\text{proj}(x_{j}^{*})$, for $j=1,2$, and $\bm{\kappa}(s_{0})=\Omega^{1/2}({\bm{X}}(0)-{\bm{x}}^{(j_{0})}(0))$. For $i=1,2,\dots$
@@
    compute $U=\text{proj}({\bm{X}}(t_{i}))$ and set $j_{i}\in\{1,2\}$ such that $u_{j_{i}}^{*}\,U>0$ and $s_{i}=\operatorname*{argmin}_{t\geq 0}\lVert U-u_{j_{i}}(t)\rVert$;
@@ 5 Numerical investigations
Next, we study the accuracy and computational efficiency of the pcLNA compared to the SSA. For this, we separately use SSA and pcLNA to produce a large number of long-time, stochastic trajectories and then compare their empirical probability distribution at various time-points. We perform this exercise for various systems, either presenting bi-stability or a Hopf bifurcation.
For Hopf bifurcation systems, we perform comparisons for all three qualitatively different cases where the system complex conjugate pair of eigenvalues in (H.II) has $\mu(\alpha)<0,=0,>0$. In these three cases, the limiting RRE system presents quickly dampened oscillations, slowly dampened oscillations, and sustained oscillations, respectively (see also Figure 1A and 4). For bistable systems, we focus on parameter values satisfying 1 (see Figure 1B, right) where the system present bistability.
@@
The pcLNA algorithm 4 requires the choice of at least two solutions of RRE that differ only by their initial condition and converge to a different equilibrium. In the presented simulations, the two initial conditions are of the form $(1-\epsilon)\,{\bm{x}}_{\alpha}+\epsilon\,{\bm{x}}_{j}^{*}+{\bm{c}}$, $j=1,2$. Here $0\leq\epsilon\leq 0.1$, ${\bm{x}}_{\alpha}$ the unstable equilibrium and ${\bm{x}}_{j}^{*}$, $j=1,2$, the stable equilibria. The real vector $\bm{c}$} performs a change of location in a direction orthogonal to the eigenvector corresponding to the real positive eigenvalue of the Jacobian $\bm{J}({\bm{x}}^{u})$. This change of location allows to view some short-time transient dynamics, but it is not necessary if one is interested in the long-time dynamics. For the specific initial conditions and other details, see SI Sections S4.4- S4.6.

### v4

@@ 4.1 The pcLNA algorithm for Hopf bifurcation systems
For a reaction network with macroscopic RRE as in (6) satisfying conditions (H.I), (H.II), we now wish to define the projection metric $d$ and set of reference trajectories $\Gamma$ to define the phase-correcting map $G$, as described in Section 3.3.4. Consider some value $\alpha$ of the bifurcation parameter that is near the critical $\alpha_{p}$. The normal form dynamics imply that trajectories starting from different initial conditions rapidly converge to a common path. This motivates the use of a single deterministic trajectory $\Gamma=\{{\bm{x}}(t):t\geq 0\}$ near ${\bm{x}}_{\alpha_{p}}$, which proved sufficient to achieve accurate simulations for all systems considered in Section 6.
To obtain a readily applicable $G$ using the steps in Section 3.3.4 we identify the conjugate pair of eigenvectors, $\hat{{\bm{u}}}_{1}\pm i\hat{{\bm{u}}}_{2}$, corresponding to the imaginary eigenvalues of the Jacobian $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ as in (H.II). Then, the eigenspace corresponding to these complex eigenvalues is spanned by $\{\hat{{\bm{u}}}_{1},\hat{{\bm{u}}}_{2}\}$. Letting ${\bm{R}}_{c}=[\hat{{\bm{u}}}_{1}\;\hat{{\bm{u}}}_{2}]$ we define the map $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}^{2}$, $\text{proj}({\bm{x}})={\bm{R}}_{c}^{-1}({\bm{x}}-{\bm{x}}_{\alpha_{p}})$. Now $d$ and $G$ can be defined as in (28), and (29) with $L=1$; phase correction occurs to minimize the distance in the approximate centre coordinates between the stochastic state ${\bm{U}}=\text{proj}({\bm{X}})$ and single deterministic path $\{{\bm{u}}(t)=\text{proj}({\bm{x}}(t)):t\geq 0\}$. Having defined the metric $d$ for the map G{G}, one may follow the steps in Algorithm 1 to produce simulations of Hopf bifurcation systems.
Since only one RRE solution $\Gamma$ is required for this type of system the map $G$ can be written as the convolution $G={\bm{x}}\circ\mathcal{M}\circ\text{proj}$ where $\mathcal{M}:\mathbb{R}^{2}\to[0,\infty)$, such that $s=\mathcal{M}({\bm{U}})=\operatorname*{argmin}_{t\geq 0}\lVert{\bm{U}}-{\bm{u}}(t)\rVert$ (see Figure 6). For ease, we provide a pcLNA algorithm specifically for Hopf bifurcation systems using this notation in Algorithm 2. In Section 6, we show that this algorithm is sufficient to achieve high levels of accuracy.
@@
a. Initial conditions: $t_{0}=0$, ${\bm{x}}(0)$, ${\bm{X}}(0)$, Parameters: $\Delta t$, $\Omega$.
b. System satisfying (H.I), (H.II), with solution $\Gamma=\{{\bm{x}}(t):t\geq 0\}$ of (6) and the corresponding solutions $\bm{C}$} and $\bm{D}$} of (12), (13), respectively.
c. The projection map $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}^{2}$, $\text{proj}({\bm{x}})={\bm{R}}_{c}^{-1}({\bm{x}}-{\bm{x}}_{\alpha_{p}})$ where ${\bm{R}}_{c}=[\hat{{\bm{u}}}_{1}\;\hat{{\bm{u}}}_{2}]$ with $\hat{{\bm{u}}}_{1}$, $\hat{{\bm{u}}}_{2}$ an eigenbasis corresponding to the conjugate pair of imaginary eigenvalues of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$.
Steps: Compute $\{{\bm{u}}(t):t\geq 0\}$, where ${\bm{u}}(t)=\text{proj}({\bm{x}}(t))$ and set $s_{0}$ such that $G({\bm{X}}(0))={\bm{x}}(s_{0})$ with $\bm{\kappa}(s_{0})=\Omega^{1/2}({\bm{X}}(0)-{\bm{x}}(s_{0}))$.
@@ 5 Bi-stable systems
We now consider systems that present bi-stability and wish to derive stochastic trajectories $\{{\bm{X}}(t_{i}):i=0,1,2,\dots\}$, with ${\bm{X}}(t_{i})\in\mathbb{R}^{n}$, for a reaction network with RRE limit (as $\Omega\to\infty$) satisfying (6) and Jacobian, $\bm{J}({\bm{x}},\alpha)$, as in (7) that satisfy the following conditions:
@@
    for $\alpha$ near $\alpha_{p}\in\mathbb{R}$ there exists ${\bm{x}}^{(1)}_{*}\in\mathbb{R}^{n}$ such that ${\bm{F}}({\bm{x}}^{(1)}_{*},\alpha)={\bm{0}}$ and all the eigenvalues of the Jacobian matrix $\bm{J}({\bm{x}}^{(1)}_{*},\alpha)$ are negative.
The last condition (B.III) ensures the presence of a stable equilibrium point for all values of $\alpha$ near $\alpha_{p}$. The conditions (B.I-II) prescribe the presence of a non-hyperbolic equilibrium, ${\bm{x}}_{\alpha_{p}}$, and the existence of a parameter dependent, one-dimensional, attracting manifold near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$ for which (6) is l.t.e. to a system
[…]
where for each $t\geq 0$, $u(t)\in\mathbb{R},\ {\bm{v}}(t)\in\mathbb{R}^{n-1},\ f:\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{2}$ is smooth and ${\bm{A}}(u(t),\alpha)$ is an $(n-1)\times(n-1)$ matrix with all of its eigenvalues having negative real part. For the generic fold bifurcation system (31), the non-hyperbolic equilibrium present at $\alpha_{p}$ disappears for $\alpha<\alpha_{p}$ while two equilibrium points, one stable and one unstable, emerge for $\alpha>\alpha_{p}$. Therefore, a system satisfying the conditions (B.I-III) presents two stable and one unstable equilibrium points for $\alpha>\alpha_{p}$. That is, we arrive to the following result.
@@
For a system as in (6) satisfying the conditions (B.I)-(B.III), with value of $\alpha$ such that $\alpha-\alpha_{p}$ is positive and sufficiently small, there exist ${\bm{x}}^{(1)}_{*},{\bm{x}}_{\alpha},{\bm{x}}^{(2)}_{*}\in\mathbb{R}^{n}$ with ${\bm{F}}({\bm{x}}^{(1)}_{*},\alpha)={\bm{F}}({\bm{x}}_{\alpha},\alpha)={\bm{F}}({\bm{x}}^{(2)}_{*},\alpha)={\bm{0}}$ and all the eigenvalues of the Jacobian matrices $\bm{J}({\bm{x}}^{(1)}_{*},\alpha)$ and $\bm{J}({\bm{x}}^{(2)}_{*},\alpha)$ are negative (linear stability), while $\bm{J}({\bm{x}}_{\alpha},\alpha)$ has one positive eigenvalue (instability) and the rest are negative.
Similarly to Hopf bifurcation systems in Section 4, the image of the LNA, ${\bm{X}}(t)={\bm{x}}(t)+\Omega^{-1/2}\bm{\xi}(t)$ under the local homeomorphism reveals stochastic processes $U(t)\approx u(t)+\Omega^{-1/2}\ \xi_{u}(t)$, evolving in the single centre coordinate; and ${\bm{V}}(t)\approx{\bm{v}}(t)+\Omega^{-1/2}\bm{\xi}_{\bm{v}}(t)$, evolving in the stable coordinates. Their stochastic perturbations, $\xi_{\bm{u}}(t)$ and $\bm{\xi}_{\bm{v}}(t)$, evolve according to (20).
@@ 5.1 The pcLNA algorithm for bi-stable systems
A system satisfying Remark 1 with $\alpha>\alpha_{p}$ exhibits two stable equilibrium points. The standard LNA, which relies on a single deterministic solution to simulate a trajectory, will produce paths centered around the “nearest” equilibrium. As a result, it will fail to approximate trajectories produced by the SSA whenever there is a substantial probability that a stochastic trajectory produced by the SSA escapes from this equilibrium (see Figure 2(D) and SI Figure S19).
Here we discuss how to choose a metric $d$ and a set of reference trajectories $\Gamma$ (following the procedure in 3.3.4) in order to apply the pcLNA algorithm in 1 and obtain accurate simulations.
If ${\bm{R}}_{c}={\bm{u}}_{1}$ is the eigenvector corresponding to the zero eigenvalue of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$, then we define the projection map $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}$, onto the eigenspace that assigns to each state $\bm{x}$} its approximate location in the centre coordinate $u=\text{proj}({\bm{x}})={\bm{R}}_{c}^{-1}({\bm{x}}-{\bm{x}}_{\alpha_{p}})$. Again, the distance metric in (28) is $d({\bm{X}},{\bm{x}})=|U-u|,$ where $U=\text{proj}({\bm{X}})$, and $u=\text{proj}({\bm{x}})$.
When selecting the set of reference trajectories $\Gamma$, at least two trajectories—each converging to a different equilibrium—are required. This is necessary to allow stochastic trajectories that start from the same initial condition to reach the vicinity of different equilibria via LNA transitions, a behavior that is observed in SSA trajectories.
Having defined the distance metric $d$ and the set of reference trajectories $\Gamma$, we can apply (15) and (16) to define the phase-correcting map $G$ and therefore apply the Algorithm 1.
With this in place, we now present a pcLNA algorithm specifically for bistable systems in Algorithm 3. In Section 6, we show that this simple phase map is sufficient to achieve high levels of accuracy.
@@
c. The projection map $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}$, $u=\text{proj}({\bm{x}})={\bm{R}}_{c}^{-1}({\bm{x}}-{\bm{x}}_{\alpha_{p}})$ where ${\bm{R}}_{c}={\bm{u}}$ with $\bm{u}$} a normalised eigenvector corresponding to the zero eigenvalue of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$.
Steps: Compute $\{u^{(j)}(t):t\geq 0\}$ where $u^{(j)}(t)=\text{proj}({\bm{x}}^{(j)}(t))$ for $j=1,2$, and $\bm{\kappa}(s_{0})=\Omega^{1/2}({\bm{X}}(0)-{\bm{x}}^{(j_{0})}(0))$. For $i=1,2,\dots$
@@
    draw sample $\bm{\xi}(s_{i-1}+\Delta t)$ from the MVN distribution with mean ${\bm{C}}^{(j_{i-1})}(s_{i-1},s_{i-1}+\Delta t)\bm{\kappa}(s_{i-1})$, and variance matrix ${\bm{D}}^{(j_{i-1})}(s_{i-1},s_{i-1}+\Delta t)$;
@@
    compute $U=\text{proj}({\bm{X}}(t_{i}))$ and find $(j_{i},s_{i})=\operatorname*{argmin}_{\ell\in\{1,\dots,L\},t\geq 0}\|U-u^{(\ell)}(t)\|$;
@@ 6 Numerical investigations
Next, we study the accuracy and computational efficiency of the pcLNA compared to the SSA. For this, we separately use SSA and pcLNA to produce a large number of long-time, stochastic trajectories and then compare their empirical probability distribution at various time-points. We perform this exercise for various systems, either presenting bi-stability or a Hopf bifurcation. All numerical investigations were performed on a MacBook Pro workstation with an Apple M2 processor (8 CPU cores) and 16 GB RAM using the software PeTTSy [10] in MATLAB. [27].
For Hopf bifurcation systems, we perform comparisons for all three qualitatively different cases where the system complex conjugate pair of eigenvalues in (H.II) has $\mu(\alpha)<0,=0,>0$. In these three cases, the limiting RRE system presents quickly dampened oscillations with a focus shape, slowly dampened oscillations, and sustained oscillations, respectively (see also Figure 1A and 4). For bistable systems, we focus on parameter values satisfying 1 (see Figure 1B, right) where the system present bistability.
@@
The pcLNA algorithm 3 requires the choice of at least two solutions of RRE that differ only by their initial condition and converge to a different equilibrium. In the presented simulations, the two initial conditions are of the form $(1-\epsilon)\,{\bm{x}}_{\alpha}+\epsilon\,{\bm{x}}^{(j)}_{*}+{\bm{c}}$, $j=1,2$. Here $0\leq\epsilon\leq 0.1$, ${\bm{x}}_{\alpha}$ the unstable equilibrium and ${\bm{x}}^{(j)}_{*}$, $j=1,2$, the stable equilibria. The real vector $\bm{c}$} performs a change of location in a direction orthogonal to the eigenvector corresponding to the real positive eigenvalue of the Jacobian $\bm{J}({\bm{x}}^{u})$. This change of location allows to view some short-time transient dynamics, but it is not necessary if one is interested in the long-time dynamics. For the specific initial conditions and other details, see SI Sections S4.4-S4.6.
@@
Finally, to quantify the computational cost of SSA and pcLNA across systems, we report the CPU times in Table 1. We also introduce a simple and practical predictor of the SSA runtime based on the system’s average reaction activity.
For a given system, the total propensity $\pi_{0}({\bm{Y}})=\sum_{j=1}^{r}\pi_{j}({\bm{Y}}),$ is evaluated along the deterministic trajectories $\{{\bm{x}}^{(\ell)}(t):t\in[0,T],\ell\in\mathcal{L}\}$ used for pcLNA simulation taking ${\bm{Y}}=\Omega\,{\bm{x}}^{(\ell)}(t)$ (here $\mathcal{L}=\{1,\dots,L\},\,L\in\{1,2\}$). This yields the time-dependent reaction activity of the mean-field dynamics which, although it neglects stochastic deviations, provides a reliable summary of the system’s reaction intensity over different dynamical phases. We then define the average total propensity
$\bar{\pi}_{0}=\int_{0}^{T}\pi_{0}(\Omega{\bm{x}}(t))\,dt.$
Since the SSA explicitly simulates every reaction event, and each event has constant computational cost, the expected number of reactions in time $T$, given by $N_{r}=T\bar{\pi}_{0}$, provides an easily-computable, interpretable predictor of SSA runtime.
[…]
CPU time
System $N_{r}/10^{3}$ $n$ $r$ SSA pcLNA
toggle-switch [14] 5.0347 2 4 0.0296 0.0110
cell-cycle [2] 4.4595 2 6 0.0318 0.0065
somitogenesis-switch [21] 821.18 4 10 4.6188 0.0690
smallest Hopf [71] (focus) 376.187 3 5 2.7065 0.0184
smallest Hopf [71] (Hopf) 468.307 3 5 3.3828 0.0181
smallest Hopf [71] (limit-cycle) 564.457 3 5 3.9284 0.0173
brusselator [39] (focus) 360.305 2 4 3.4655 0.0196
brusselator [39] (Hopf) 405.078 2 4 3.8119 0.0171
brusselator [39] (limit-cycle) 550.039 2 4 3.2390 0.0417
[…]
Table 1: Computational time of SSA and pcLNA for different systems. The median CPU times for the SSA and pcLNA simulations with parameter values given in the SI Sections S4.1-S4.6 are displayed along with number of populations ($n$), reaction channels ($r$), and average number of reaction occurrences in thousands ($N_{r}$, see 6).